31 August 2018 MST - Night 7 - Install 20 iter FF from 30th, IF Data (20180901 UT)

Summary

Date: Sat, 1 Sep 2018 06:18:17 -0700
From: Doug Miller 
To: AOGroup 

This is the last night we will perform 4D DX AdSec calibration.  Tonight,
Saturday Night, the telescope opens for the 2018b Observer Semester.

Tonight we were able to collect a couple sets of Influence Function data
with the LUCI Cryopumps off.  However, it looks like there is some
transient vibration source, possibly the Summer Chiller, so we took a partial
set of data with the Summer Chiller Off.

I will have to analyze the data to determine if we have enough good data
for a complete set of IF data.  Hopefully later today I will have the noise
data sets reduced so we can try to link the transient vibration to telescope
systems turning on and off.

Thanks to Steve and Josh for operation the telescope in a non-standard
way and turning the LUCI Cryopumps off and on.

Thanks to Dave T. for waking up every couple hours through the night
to check on the LUCIs.  It sounds like we did not warm them up too much
and they should be back to normal temps in a couple hours.

Plot Descriptions

On this night we collect three sets of Influence Function data. The basic setup of the telescope was:
  • M1s were down
  • Tree house fans were On
  • LUCI Cryopumps were off

The Sets are
  1. IF_1 Full Set (0-649 modes)
  2. IF_2 Partial Set (0-433 modes)
  3. IF_s Partial set (0-181 modes)

The motivation for each set was
  1. We thought that the LUCI Cryopumps were the major contributor to vibrations seen by the 4D. Thus with them off, the data should be good
    • After analyzing the vibration data (sets with _noise in the name) we saw that at 08:26:00 the noise decreased dramatically.
  2. Thinking that the Summer Chiller may have cycled off just before 08:26 we decided to collect another set of IF data. Modes 434-649 from IF_1 set had very low noise so only modes 0-433 were collect.
  3. We still had time remaining that the LUCI Cryopumps could be off so we forced the Summer Chiller to be off and collected data until the HBS oil nearly reached Alarm state. This allowed us to collect modes 0-181.

Results I think we can combine modes 434-649 from IF_1 set and modes 0-433 from IF_2 set and have a full set of low noise data. TBD

The each plot has
  • The "Good" data from 20130826 that had low vibration and resulted in a good KL basis set (black line)
  • The vibration data of the newly changed state of the telescope (i.e. the vibration plots after the listed system was turned off) (red line)
  • The vibration data of the previous state of the telescope, for comparison (green or blue line, see Note below)

For each vibration data set there are three plots.
  1. The RMS Tip/Tilt noise as a function of the number of samples averaged. For the KL basis set calculation we average 125 samples for low order modes (0-9), 51 samples for middle modes (10-20) and 35 samples for high modes (21-649)
  2. The full vibration spectrum is calculated with 1000 samples. The 4D sample rate is 28 Hz so the plotted spectrum is from 0 to 14 Hz.
  3. An expanded section of the vibration spectrum from 10-14 Hz. The Y scale is the same as the previous plot.
Note The line colors are not consistent from the noise plot to the vibration spectrum plots. In the noise plot the Green line is the new state of the telescope and in the vibration spectra it is the Blue line.

Note If you wish to see an enlarge version of the plot scroll to the bottom of the page and click the name of the image in the table.

Definitions (don't worry about it)
  • Set on = led-cup-green-plus
  • Set off = led-cup-red-close
Note In the tables below the system state is
  • On led-cup-green-plus
  • Off led-cup-red-close

led-green Install 20 Iter FF from 20180830

05:00 Install 20 iter FF Matrix
  • /local/aomeas/adsec_calib/ff_matrix/20180830_120714/
  • aoroot
  • confcalib

led-green Flatten the Mirror

05:38 Flatten the mirror
  • Apply 400 mode recon, 32um
  • hexalign()
  • Apply 600 mode recon, 24 nm
    • /local/aomeas/adsec_calib/flat/20180901_053816/flattening_result.sav
    • Flat with 600 modes:
      flat600_1_1.jpg
    • Flat with 600 modes:
      flat600_1_2.jpg
    • Flat with 600 modes:
      flat600_1_3.jpg
  • Apply 640 mode recon, 24 nm
    • /local/aomeas/adsec_calib/flat/20180901_055206/flattening_result.sav
  • hexalign()
    • Flat with 640 modes:
      flat640_1_1.jpg
    • Flat with 640 modes:
      flat640_1_2.jpg
    • Flat with 640 modes:
      flat640_1_3.jpg

led-green Vibration Sets

Times are UT on 20180901 UT.

05:56 Vibration Set
  • 20180901_055600_noise
  • 20180901_055600_noise

06:00 Steve bring Primaries down
  • Dave turn MCE/ROE off on both LUCIs
  • A lot of vibrations
  • 1_Primary_Down.mov: Movie of 4D image with Everything running but Primary

06:41 Shut off some Systems
  • Az and El Drive Off
  • Enc drives are off
  • HBS Off

06:42 Turn off summer chiller
  • 4D looks better

07:00 Flatten with 600 modes
  • save flat
  • Set at default.sav
  • apply flat, 376 jumps in curr
  • flat_w_376_cur.jpg
  • apply old flat and still bump
  • Rest and Set: 646 modes. No bump

07:20 Flatten with 600 modes
  • /local/aomeas/adsec_calib/flat/20180901_072057/flattening_result.sav

led-green Collect IF Data, Set 1

07:40 Some things shut off
  • Primaries Rested
  • Both LUCI Cold Heads Off
  • Summer Chiller ON (Needed for HBS)
  • HBS On

07:34 Cold Head Off

07
  • print, i4d_opt_intmat(0, 4, AMPRMS=500e-9, n=125)

Mode Amplitude Samples Tracking Number
0-4 500nm 125 20180901_074327
0-4 1000nm 125 20180901_074445
20180901_074500_noise
5-9 350 nm 125 20180901_074648
5-9 7000 nm 125 20180901_074736
10-19 250nm 51 20180901_074844
10-19 500nm 51 20180901_074925
20180901_075000_noise + hexalign()
20-37 300 35 20180901_075058
20-37 150 35 20180901_075152
38-55 300 35 20180901_075246
38-55 150 35 20180901_075328
56-73 250 35 20180901_075430
56-73 150 35 20180901_075519
20180901_075500_noise + hexalign()
74-91 150 35 20180901_075838
92-109 100 35 20180901_075943
110-127 100 35 20180901_080035
128-145 80 35 20180901_080133
146-163 80 35 20180901_080222
164-181 80 35 20180901_080328
182-199 80 35 20180901_080420
20180901_080500_noise
200-217 70 35 20180901_080713
218-235 50 35 20180901_080852
236-253 50 35 20180901_080958
254-271 50 35 20180901_081156
272-289 50 35 20180901_081350
290-307 40 35 20180901_081444
308-325 40 35 20180901_081535
20180901_081600_noise
326-343 40 35 20180901_081914
344-361 40 35 20180901_082019
362-370 40 35 20180901_082128
371-379 40 35 20180901_082235
380-397 40 35 20180901_082323
398-415 40 35 20180901_082423
416-433 40 35 20180901_082524
20180901_082600_noise
434-451 40 35 20180901_082811
452-469 40 35 20180901_082919
470-487 40 35 20180901_083014
488-505 40 35 20180901_083115
506-523 40 35 20180901_083213
524-541 40 35 20180901_083311
20180901_083400_noise
542-559 40 35 20180901_083550
560-577 40 35 20180901_083658
578-595 40 35 20180901_083758
596-613 40 35 20180901_083856
614-631 40 35 20180901_084002
632-649 40 35 20180901_084136
20180901_084200_noise

  • All Plots IF_1:
    all_if_1.jpg
  • All Plots IF_1 with low noise after 08:26:
    all_if_1_2.jpg

Tree House Fans M1s Lower HBS Pump #2 LUCI1 MCE/ROE LUCI1 Coldhead LUCI1 Rack LUCI2 MCE/ROE LUCI2 Coldheads LUCI2 Rack ULTH Fans Summer Chiller Glycol Circ Pumps ICS Stow Pins In HBS On M3 Power M2 Hex Power Air Comp Acopian
led-cup-green-plus led-cup-red-close led-cup-green-plus led-cup-red-close led-cup-red-close led-cup-green-plus led-cup-red-close led-cup-red-close led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus

High Vibration Movie

  • Note All frames are smooth
  • These data were collected with the comm4d.capture()
  • 081600.mov: High Vibration 081600

Low Vibration Movie

  • Note Every 10th frame are "bad" images with jumps and pixilation.
  • These data were collected with the comm4d.capture()
  • After this noise data set all Influence Function and Noise data have the same occasional "bad" image
  • 082600.mov: Low Vibration 082600

led-green Collect IF Data, Set 2

11:15 LUCI Cryopumps Off

Mode Amplitude Samples Tracking Number
20180901_112100_noise
0-4 500nm 125 20180901_112339
0-4 1000nm 125 20180901_112431
5-9 350 nm 125 20180901_112523
5-9 700 nm 125 20180901_112621
10-19 250nm 51 20180901_112718
10-19 500nm 51 20180901_112815
20180901_112900_noise
20-37 300 35 20180901_113322
20-37 150 35 20180901_113457
38-55 300 35 20180901_113709
38-55 150 35 20180901_113847
56-73 250 35 20180901_113950
56-73 150 35 20180901_114101
20180901_114100_noise
74-91 150 35 20180901_114325
92-109 100 35 20180901_114422
110-127 100 35 20180901_114512
128-145 80 35 20180901_114607
146-163 80 35 20180901_114714
164-181 80 35 20180901_114831
182-199 80 35 20180901_115021
20180901_115100_noise
200-217 70 35 20180901_115403
218-235 50 35 20180901_115518
236-253 50 35 20180901_115616
254-271 50 35 20180901_115718
272-289 50 35 20180901_115810
290-307 40 35 20180901_115910
308-325 40 35 20180901_120010
20180901_120000_noise
326-343 40 35 20180901_120250
344-361 40 35 20180901_120342
362-370 40 35 20180901_120441
371-379 40 35 20180901_120810
380-397 40 35 20180901_120853
398-415 40 35 20180901_120955
416-433 40 35 20180901_121052
416-433 40 35 20180901_121151
20180901_121300_noise

  • All Plots IF 2:
    all_if_2.jpg

Tree House Fans M1s Lower HBS Pump #2 LUCI1 MCE/ROE LUCI1 Coldhead LUCI1 Rack LUCI2 MCE/ROE LUCI2 Coldheads LUCI2 Rack ULTH Fans Summer Chiller Glycol Circ Pumps ICS Stow Pins In HBS On M3 Power M2 Hex Power Air Comp Acopian
led-cup-green-plus led-cup-red-close led-cup-green-plus led-cup-red-close led-cup-red-close led-cup-green-plus led-cup-red-close led-cup-red-close led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus

led-green Collect IF Data, Set 3 with Summer Chiller Off

12:15 Josh disable Summer Chiller. Reports that it is off

Mode Amplitude Samples Tracking Number
20180901_121500_noise
0-4 500nm 125 20180901_121722
5-9 350 nm 125 20180901_121823
10-19 250nm 51 20180901_121918
10-19 250nm 51 20180901_122005
10-19 250nm 51 20180901_122059
20-37 150 35 20180901_122143
38-55 150 35 20180901_122239
56-73 150 35 20180901_122333
74-91 150 35 20180901_122424
20180901_122500_noise
92-109 100 35 20180901_122619
110-127 100 35 20180901_122711
128-145 80 35 20180901_122806
146-163 80 35 20180901_122857
164-181 80 35 20180901_122946
20180901_123000_noise
HBS temperature increase 8-9 C. Alarm at 10. Stop HBS and turn on Cooling

  • All Plots IF 3:
    all_if_3.jpg

Tree House Fans M1s Lower HBS Pump #2 LUCI1 MCE/ROE LUCI1 Coldhead LUCI1 Rack LUCI2 MCE/ROE LUCI2 Coldheads LUCI2 Rack ULTH Fans Summer Chiller Glycol Circ Pumps ICS Stow Pins In HBS On M3 Power M2 Hex Power Air Comp Acopian
led-cup-green-plus led-cup-red-close led-cup-green-plus led-cup-red-close led-cup-red-close led-cup-green-plus led-cup-red-close led-cup-red-close led-cup-green-plus led-cup-green-plus led-cup-red-close led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus led-cup-green-plus

-- %USERSIG{DougMiller - 2018-09-01}%

  • Chiller Temperatures from LBTplot - summer chiller is running when green line goes down.:
    Screen Shot 2018-09-01 at 11.48.32 AM.png
I Attachment Action Size Date Who Comment
081600.movmov 081600.mov manage 19 MB 06 Sep 2018 - 16:29 DougMiller High Vibrations 081600
082600.movmov 082600.mov manage 14 MB 06 Sep 2018 - 16:30 DougMiller Low Vibrations 082600
1_Primary_Down.movmov 1_Primary_Down.mov manage 1 MB 01 Sep 2018 - 06:22 DougMiller Movie of 4D image with Everything running but Primary
Screen Shot 2018-09-01 at 11.48.32 AM.pngpng Screen Shot 2018-09-01 at 11.48.32 AM.png manage 494 K 12 Oct 2018 - 23:34 JohnHill Chiller Temperatures from LBTplot - summer chiller is running when green line goes down.
all_if_1.jpgjpg all_if_1.jpg manage 132 K 04 Sep 2018 - 21:03 DougMiller All Plots
all_if_1_2.jpgjpg all_if_1_2.jpg manage 95 K 04 Sep 2018 - 21:03 DougMiller All Plots
all_if_2.jpgjpg all_if_2.jpg manage 142 K 04 Sep 2018 - 09:12 DougMiller All Plots
all_if_3.jpgjpg all_if_3.jpg manage 105 K 04 Sep 2018 - 09:12 DougMiller All Plots
flat600_1_1.jpgjpg flat600_1_1.jpg manage 173 K 01 Sep 2018 - 06:20 DougMiller Flat with 600 modes
flat600_1_2.jpgjpg flat600_1_2.jpg manage 102 K 01 Sep 2018 - 06:20 DougMiller Flat with 600 modes
flat600_1_3.jpgjpg flat600_1_3.jpg manage 201 K 01 Sep 2018 - 06:20 DougMiller Flat with 600 modes
flat640_1_1.jpgjpg flat640_1_1.jpg manage 172 K 01 Sep 2018 - 06:21 DougMiller Flat with 640 modes
flat640_1_2.jpgjpg flat640_1_2.jpg manage 98 K 01 Sep 2018 - 06:21 DougMiller Flat with 640 modes
flat640_1_3.jpgjpg flat640_1_3.jpg manage 205 K 01 Sep 2018 - 06:21 DougMiller Flat with 640 modes
Topic revision: r12 - 13 Oct 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback